# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

### v5

@@ 1.1 Background
Randomized controlled trials (RCTs) are widely regarded as the gold standard for evaluating the effects of treatments or interventions (Neyman, 1923; Rubin, 1974). Through randomization, covariates are balanced on average across treatment groups, enabling unbiased estimation of treatment effects. However, in practice, randomization does not guarantee perfect balance in finite samples, particularly when dealing with a large number of covariates. Such covariate imbalances can obscure the interpretation of trial outcomes and introduce accidental biases. Incorporating covariates into the analysis can enhance both the interpretability and efficiency of RCT results. While covariate balance can be pursued during both the design (Fisher, 1926; Taves, 1974; Pocock and Simon, 1975; Rosenberger and Sverdlov, 2008; Hu and Hu, 2012; Hu et al., 2014; Morgan and Rubin, 2012; Li et al., 2018b; Zhao and Ding, 2024) and analysis phases of an RCT, this paper focuses on methods for covariate adjustment during the analysis of already-conducted trials.
@@ 1.2 Related work
Several methods have been studied for data-driven variable selection procedures in RCTs. For example, the least absolute shrinkage and selection operator (Lasso) (Tibshirani, 1996) has been used in conjunction with ANHECOVA to select relevant covariates (Bloniarz et al., 2016; Schochet, 2022; Liu et al., 2023; Zhu et al., 2024). Another approach involves selecting the top $k$ covariates with the highest marginal correlations with the outcome, where $k<p$ (Senn, 1994). Alternatively, covariates can be screened using preliminary significance tests, such as two-sample $t$-tests or $z$-tests, to identify those most strongly associated with the treatment assignment or outcome (Permutt, 1990; Zhao and Ding, 2023). These methods have primarily focused on continuous outcomes and the application of ANCOVA and ANHECOVA estimators. As we noted, identifying which covariate sets are most beneficial for covariate adjustment is the objective. Cho and Yang, along with additional references, propose various strategies tailored to specific purposes of covariate adjustment in observational studies. Similarly, Yang et al. (2020) offered aligned strategies in the context of data integration. In the context of RCTs, where bias is controlled, covariate adjustment primarily aims to enhance efficiency. Therefore, consistent with previous research (Cho and Yang, 2023; Yang et al., 2020), we recommend selecting covariates that are strong prognostic factors for the outcome.
@@ 1.3 Our contributions
Variable selection Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
@@ 3.2 Theoretical results
We establish theoretical results for efficiency gains via the AIPW estimator, with detailed proofs provided in Web Appendix A of our Online Supplemental Material. The key results are summarized in the following Theorem 1.
@@
For point (ii), conditions (a) and (b) have the following interpretations: for (a), the GLM must correctly capture the sign of the true conditional outcome with probability 1; and for (b), the bias of the GLM is constrained to maintain efficiency. For example, with positive-valued outcomes, conditions (a) and (b) can be simplified to $0\leq g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})\leq 2\mathbb{E}(Y(a)\mid\mathbf{X})$, which is reasonable for some positive categorical outcomes, such as binary outcomes modeled by logistic regression. Further illustration and discussion of these conditions are provided in Web Appendix A of our Online Supplemental Material.
@@ 3.3 Post-selection inference
Finally, we remark on the robustness of variance estimators for covariate-adjusted estimators. Thanks to the RCT design, the propensity score (treatment model) is known and correctly specified. In our framework, variable selection is applied only to the outcome models used in the covariate-adjusted estimators. For all estimators considered—ANCOVA, ANHECOVA, and AIPW with Lasso-based variable selection—Condition 1 ensures that the residuals in their HW sandwich or model-based variance estimators capture the unexplained variation, including uncertainty due to variable selection (Liu et al., 2023; Bloniarz et al., 2016; Zhu et al., 2024). As a result, the robust variance estimators constructed from these residuals implicitly account for this uncertainty.
@@ 4.1 Data generating process
for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect. The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material (Section B) for reproducibility.
@@ 4.2 Methods for comparison
(Senn, 1994; Zhao and Ding, 2023)
@@ 4.4 Results
For brevity, we present the simulation results for the nonlinear continuous and nonlinear outcomes under sample sizes of 40 and 200 in this section. Complete simulation results, including figures, are available in our Online Supplemental Material (Section B).
@@
In terms of CP% and power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size. However, even when $N=200$, we observe below nominal or missing CP% for all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations. Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CP% under the continuous outcome with $N=200$. However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material. This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research. Overall, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well (in terms of both point and variance estimations) across both sample sizes and both outcomes.
@@
Finally, to empirically assess the asymptotic normality of post-selection estimators following Lasso and adaptive Lasso, we present histograms of AIPW estimates after variable selection in Figure 4. Additional results for ANCOVA, ANHECOVA, and other simulation scenarios are provided in the Online Supplemental Material. As shown in Figure 4, the empirical distributions closely align with the corresponding theoretical normal densities (indicated by red dashed lines). As sample size increases, the estimates become more concentrated around the mean, and the agreement with the theoretical curves improves under both continuous and binary outcomes. Supplemental figures further support these findings, demonstrating strong normality even at moderate sample sizes (e.g., $N=100$ and $N=500$). Taken together, these results provide empirical evidence for the asymptotic normality of covariate-adjusted AIPW estimators following variable selection via Lasso or adaptive Lasso.
@@ 6.2 Discussion
There are also several promising avenues for extending our framework. For instance, alternative adjustment methods could be incorporated to further improve efficiency. Existing literature suggests that integrating propensity score modeling, $e(\mathbf{X})=P(A=1\mid\mathbf{X})$, into the AIPW estimator (Gao et al., 2024b), or applying propensity score weighting methods without outcome modeling (such as inverse probability weights (Shen et al., 2014), overlap weights (Li et al., 2018a; Yang et al., 2021; Zeng et al., 2021), or other schemes (Li and Greene, 2013; Liu et al., 2024b)) can enhance efficiency, even when the propensity score is known (e.g., $e(\mathbf{X})=0.5$ for 1:1 randomization across all $\mathbf{X}$) (Williamson et al., 2014; Shen et al., 2014). However, extending our framework to include propensity score modeling introduces an additional variable selection challenge, effectively leading to a double selection problem for the AIPW estimator (Koch et al., 2018; Belloni et al., 2014). Moreover, although logistic regression is commonly used for propensity score estimation (Thomas et al., 2020; Hirano and Imbens, 2001; Connors et al., 1996) it remains uncertain whether it consistently enhances efficiency or if alternative methods with clearer regularity conditions might be preferable. Another potential extension involves prognostic covariate adjustment (PROCOVA), which employs machine learning techniques to generate prognostic scores based on baseline characteristics (Harrell Jr et al., 1984, 1996; Vanderbeek et al., 2022; Rubin and Thomas, 2000; Schuler et al., 2021; Li et al., 2024). Future work could extend our variable selection methods to PROCOVA or explore the joint use of propensity and prognostic scores for covariate adjustment (Yang and Zhang, 2023; Zhang et al., 2022).
@@
Furthermore, our framework can be extended to a wide range of more general and complex settings. These include more general covariate-adaptive randomization (Bannick et al., 2025; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2024; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2024; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020). Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2024; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2023; Wang et al., 2025a). Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Banik and Shao, 2022; Wang et al., 2025b; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2025; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
@@
Acknowledgments
The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor. Yi Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896. This project is supported by the Food and Drug Administration (FDA) of the U.S. Department of Health and Human Services (HHS) as part of a financial assistance award U01FD007934 totaling \$1,674,013 over two years funded by FDA/HHS. It is also supported by the National Science Foundation under Award Number SES 2242776, totaling \$225,000 over three years. The contents are those of the authors and do not necessarily represent the official views of, nor an endorsement by, FDA/HHS, the NIH, or the U.S. Government.
@@ Acknowledgments
Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
@@ X The covariate vector/matrix/data frame $\checkmark$
lasso.family Family used in Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial" (Same as the glmnet package (Friedman et al., 2021))
@@
Both trt.name and ctrl.name need to be specified without default, which are the value of treatment and control groups in A, respectively. Note that A can have more than two levels, but we require users to input the treated (trt.name) and control (ctrl.name) each time for a pair-wise comparison. The conti.out must be specified from TRUE and FALSE as well, which allows the program to identify whether it needs to be processed as a categorical outcome for valid missing data imputation purpose. The var.sel.method argument is for the variable selection method, and we specify "Lasso" for its value. We also need to specify the model family via lasso.family if "Lasso" is used. The default is "gaussian" for using linear model with Gaussian random error assumption in Lasso regression, but users can choose other families, such as "binomial" for binary outcome if using logistic regression, and any families provided in the glmnet package.Friedman et al. (2021); Hastie et al. (2021) The arguments out0.model.aipw and out1.model.aipw are for specifying the link functions used in outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, respectively, in the AIPW estimator. For example, in the binary outcome case, we use "logit" to specify logistic regression models for both potential outcomes.
@@
When $g_{a}(\mathbf{X})=\mathbb{E}(Y(a)\mid\mathbf{X})$, representing the fully specified model for $Y(a)$, $a=0,1$, the estimator $\widehat{\tau}^{\mathcal{G}}$ achieves the semiparametric efficiency bound (Hahn, 1998; Hirano et al., 2003). However, since the true model of $Y(a)$ as a function of $\mathbf{X}$ is unknown, achieving this bound is generally not expected. Nevertheless, we can still improve efficiency through “appropriate augmentation,” even without precise model specification, compared to $\widehat{\tau}_{\text{simple}}$. This is our objective in leveraging covariate information to gain efficiency (Bannick et al., 2023).

### v6

@@ 1.1 Background
Randomized controlled trials (RCTs) are widely regarded as the gold standard for evaluating the effects of treatments or interventions (Neyman, 1923; Rubin, 1974). Through randomization, covariates are balanced on average across treatment groups, enabling unbiased estimation of treatment effects. However, in practice, randomization does not guarantee perfect balance in finite samples, particularly when dealing with a large number of covariates. Such covariate imbalances can obscure the interpretation of trial outcomes and introduce accidental biases. Incorporating covariates into the analysis can enhance both the interpretability and efficiency of RCT results. While covariate balance can be pursued during both the design (Fisher, 1926; Taves, 1974; Pocock and Simon, 1975; Rosenberger and Sverdlov, 2008; Hu and Hu, 2012; Hu et al., 2014; Morgan and Rubin, 2012; Li et al., 2018b; Zhao and Ding, 2024b) and analysis phases of an RCT, this paper focuses on methods for covariate adjustment during the analysis of already-conducted trials.
@@ 1.2 Related work
Several methods have been studied for data-driven variable selection procedures in RCTs. For example, the least absolute shrinkage and selection operator (Lasso) (Tibshirani, 1996) has been used in conjunction with ANHECOVA to select relevant covariates (Bloniarz et al., 2016; Schochet, 2022; Liu et al., 2023; Zhu et al., 2025). Another approach involves selecting the top $k$ covariates with the highest marginal correlations with the outcome, where $k<p$ (Senn, 1994). Alternatively, covariates can be screened using preliminary significance tests, such as two-sample $t$-tests or $z$-tests, to identify those most strongly associated with the treatment assignment or outcome (Permutt, 1990; Zhao and Ding, 2024c). These methods have primarily focused on continuous outcomes and the application of ANCOVA and ANHECOVA estimators. As we noted, identifying which covariate sets are most beneficial for covariate adjustment is the objective. Cho and Yang, along with additional references, propose various strategies tailored to specific purposes of covariate adjustment in observational studies. Similarly, Yang et al. (2020) offered aligned strategies in the context of data integration. In the context of RCTs, where bias is controlled, covariate adjustment primarily aims to enhance efficiency. Therefore, consistent with previous research (Cho and Yang, 2024; Yang et al., 2020), we recommend selecting covariates that are strong prognostic factors for the outcome.
@@ 1.3 Our contributions
Variable selection Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2024c)
@@ 3.2 Theoretical results
We establish theoretical results for efficiency gains via the AIPW estimator, with detailed proofs provided in Web Appendix A of Online Supplemental Material. The key results are summarized in the following Theorem 1.
@@
For point (ii), conditions (a) and (b) have the following interpretations: for (a), the GLM must correctly capture the sign of the true conditional outcome with probability 1; and for (b), the bias of the GLM is constrained to maintain efficiency. For example, with positive-valued outcomes, conditions (a) and (b) can be simplified to $0\leq g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a})\leq 2\mathbb{E}(Y(a)\mid\mathbf{X})$, which is reasonable for some positive categorical outcomes, such as binary outcomes modeled by logistic regression. Further illustration and discussion of these conditions are provided in Web Appendix A of Online Supplemental Material.
@@ 3.3 Post-selection inference
Finally, we remark on the robustness of variance estimators for covariate-adjusted estimators. Thanks to the RCT design, the propensity score (treatment model) is known and correctly specified. In our framework, variable selection is applied only to the outcome models used in the covariate-adjusted estimators. For all estimators considered—ANCOVA, ANHECOVA, and AIPW with Lasso-based variable selection—Condition 1 ensures that the residuals in their HW sandwich or model-based variance estimators capture the unexplained variation, including uncertainty due to variable selection (Liu et al., 2023; Bloniarz et al., 2016; Zhu et al., 2025). As a result, the robust variance estimators constructed from these residuals implicitly account for this uncertainty.
@@ 4.1 Data generating process
for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect. The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in Web Appendix B of Online Supplemental Material for reproducibility.
@@ 4.2 Methods for comparison
(Senn, 1994; Zhao and Ding, 2024c)
@@ 4.4 Results
For brevity, we present the simulation results for the nonlinear continuous and nonlinear outcomes under sample sizes of 40 and 200 in this section. Complete simulation results, including figures, are available in Web Appendix B of Online Supplemental Material.
@@
In terms of CP% and power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size. However, even when $N=200$, we observe below nominal or missing CP% for all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2024c), where they advise against its use in practice based on theory and simulations. Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CP% under the continuous outcome with $N=200$. However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in Web Appendix B of Online Supplemental Material. This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research. Overall, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well (in terms of both point and variance estimations) across both sample sizes and both outcomes.
@@
Finally, to empirically assess the asymptotic normality of post-selection estimators following Lasso and adaptive Lasso, we present histograms of AIPW estimates after variable selection in Figure 4. Additional results for ANCOVA, ANHECOVA, and other simulation scenarios are provided in Web Appendix B of Online Supplemental Material. As shown in Figure 4, the empirical distributions closely align with the corresponding theoretical normal densities (indicated by red dashed lines). As sample size increases, the estimates become more concentrated around the mean, and the agreement with the theoretical curves improves under both continuous and binary outcomes. Supplemental figures further support these findings, demonstrating strong normality even at moderate sample sizes (e.g., $N=100$ and $N=500$). Taken together, these results provide empirical evidence for the asymptotic normality of covariate-adjusted AIPW estimators following variable selection via Lasso or adaptive Lasso.
@@ 6.2 Discussion
There are also several promising avenues for extending our framework. For instance, alternative adjustment methods could be incorporated to further improve efficiency. Existing literature suggests that integrating propensity score modeling, $e(\mathbf{X})=P(A=1\mid\mathbf{X})$, into the AIPW estimator (Gao et al., 2024b), or applying propensity score weighting methods without outcome modeling (such as inverse probability weights (Shen et al., 2014), overlap weights (Li et al., 2018a; Yang et al., 2021; Zeng et al., 2021), or other schemes (Li and Greene, 2013; Liu et al., 2024b)) can enhance efficiency, even when the propensity score is known (e.g., $e(\mathbf{X})=0.5$ for 1:1 randomization across all $\mathbf{X}$) (Williamson et al., 2014; Shen et al., 2014). However, extending our framework to include propensity score modeling introduces an additional variable selection challenge, effectively leading to a double selection problem for the AIPW estimator (Koch et al., 2018; Belloni et al., 2014). Moreover, although logistic regression is commonly used for propensity score estimation (Thomas et al., 2020; Hirano and Imbens, 2001; Connors et al., 1996) it remains uncertain whether it consistently enhances efficiency or if alternative methods with clearer regularity conditions might be preferable. Another potential extension involves prognostic covariate adjustment (PROCOVA), which employs machine learning techniques to generate prognostic scores based on baseline characteristics (Harrell Jr et al., 1984, 1996; Vanderbeek et al., 2022; Rubin and Thomas, 2000; Li et al., 2024). Future work could extend our variable selection methods to PROCOVA or explore the joint use of propensity and prognostic scores for covariate adjustment (Yang and Zhang, 2023; Zhang et al., 2022).
@@
Furthermore, our framework can be extended to a wide range of more general and complex settings. These include more general covariate-adaptive randomization (Bannick et al., 2025; Kernan et al., 1999; Broglio, 2018; Zhu et al., 2025; Yu et al., 2024), rerandomization (Zhu and Liu, 2023; Morgan and Rubin, 2012), time-to-event outcomes (Hua et al., 2024; Yi et al., 2020; Benkeser et al., 2021), restricted mean survival times (Han, 2023; Zhong et al., 2022; Karrison and Kocherginsky, 2018; Hanada et al., 2024), cluster-randomized trials (Murray, 1998; Donner et al., 2000; Stephens et al., 2012; Zhou et al., 2022; Balzer et al., 2023), and ordinal or composite outcomes with novel estimands such as the win ratio or generalized odds ratio (Agresti, 1980; Pocock et al., 2012; Barnhart et al., 2025; Matsouaka, 2022; Oakes, 2016; Zhang and Wu, 2024), as well as the Mann-Whitney rank-based statistic (Follmann et al., 2020). Other promising directions include conformal inference on individual or heterogeneous treatment effects (Yang et al., 2024; Wang et al., 2024; Gao et al., 2024a), the integration of multi-source RCT and observational data (Zhuang et al., 2025; Lee et al., 2024b, 2023; Yang et al., 2023; Han et al., 2025; Liu et al., 2024a; Yang and Wang, 2022; Liu et al., 2025), synthetic controls (Abadie et al., 2015, 2010; Ben-Michael et al., 2021), and external controls (Li et al., 2023; Lee et al., 2024c; Gao et al., 2025; Wang et al., 2025a). Additionally, alternative strategies for estimating regression coefficients, such as debiased machine learning algorithms (Chernozhukov et al., 2018; Wang et al., 2025b; Carlsson et al., 2014), debiased regressions (Lu et al., 2023; Chang et al., 2024; Chiang et al., 2023), and linear calibration (Bannick et al., 2025; Guo and Basse, 2023; Cohen and Fogarty, 2024) are of interest for future investigation.
Acknowledgment
The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor.
Funding
Yi Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896. This project is supported by the Food and Drug Administration (FDA) of the U.S. Department of Health and Human Services (HHS) as part of a financial assistance award U01FD007934 totaling \$1,674,013 over two years funded by FDA/HHS. It is also supported by the National Science Foundation under Award Number SES 2242776, totaling \$225,000 over three years. The contents are those of the authors and do not necessarily represent the official views of, nor an endorsement by, FDA/HHS, the NIH, or the U.S. Government.
@@ Acknowledgment
Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2024a); Song et al. (2021)) $\checkmark$ $\times$ MAR
@@ X The covariate vector/matrix/data frame $\checkmark$
lasso.family Family used in Lasso regression $\times$ "gaussian" "gaussian", "binomial", "poisson", "multinomial" (Same as the glmnet package)
@@
Both trt.name and ctrl.name need to be specified without default, which are the value of treatment and control groups in A, respectively. Note that A can have more than two levels, but we require users to input the treated (trt.name) and control (ctrl.name) each time for a pair-wise comparison. The conti.out must be specified from TRUE and FALSE as well, which allows the program to identify whether it needs to be processed as a categorical outcome for valid missing data imputation purpose. The var.sel.method argument is for the variable selection method, and we specify "Lasso" for its value. We also need to specify the model family via lasso.family if "Lasso" is used. The default is "gaussian" for using linear model with Gaussian random error assumption in Lasso regression, but users can choose other families, such as "binomial" for binary outcome if using logistic regression, and any families provided in the glmnet package (Friedman et al., 2021; Hastie et al., 2021). The arguments out0.model.aipw and out1.model.aipw are for specifying the link functions used in outcome models $\mu_{0}(\mathbf{X})$ and $\mu_{1}(\mathbf{X})$, respectively, in the AIPW estimator. For example, in the binary outcome case, we use "logit" to specify logistic regression models for both potential outcomes.
@@
When $g_{a}(\mathbf{X})=\mathbb{E}(Y(a)\mid\mathbf{X})$, representing the fully specified model for $Y(a)$, $a=0,1$, the estimator $\widehat{\tau}^{\mathcal{G}}$ achieves the semiparametric efficiency bound (Hahn, 1998; Hirano et al., 2003). However, since the true model of $Y(a)$ as a function of $\mathbf{X}$ is unknown, achieving this bound is generally not expected. Nevertheless, we can still improve efficiency through “appropriate augmentation,” even without precise model specification, compared to $\widehat{\tau}_{\text{simple}}$. This is our objective in leveraging covariate information to gain efficiency (Bannick et al., 2025).